OpenStudy (anonymous):

Solve: x^2 + 11x = 12

jimthompson5910 (jim_thompson5910):

x^2 + 11x = 12 x^2 + 11x - 12 = 0 (x - 1)(x + 12) = 0 ... note: -1 + 12 = 11 and -1*12 = -12

jimthompson5910 (jim_thompson5910):

I'll let you finish up

OpenStudy (anonymous):

from @jim_thompson5910 it says (x-1)(x+12)=0 you then try, x-1=0 and x+12=0